# Quantized tensor networks for solving the Vlasov-Maxwell equations

### v3

@@ 4.3.1 Advection with time varying electric field
We perform time evolution using a second-order semi-Lagrangian (SL) scheme, fourth-order Runge-Kutta (RK4), and TDVP. The calculations use $2^{L}$ grid points along the $v_{x}$ and $v_{y}$ axes (with bounds $[-12v_{th},12v_{th})$). The test electron distribution is initialized with drift velocities $v_{x,0}=v_{y,0}=0$ and thermal velocity of 1, and is evolved until a time of 100 using time step $\Delta t$. The calculations were performed using a sequential layout with data along both axes encoded via a binary mapping. Both second-order and fourth-order centered finite difference stencils are considered.
Figure 2: Rms errors in the dynamics for advection with time-varying electric field. Test parameters are set to be $\omega=0.4567$, $E_{0}=0.9$, and $B_{0,z}=1$. Calculations are performed with bond dimension $D$, grid resolution $2^{L}$ along each dimension, and time step $\Delta t$. (top) Results for varied $D$ and fixed $L=8$ and $\Delta t=0.025$. (middle) Results for varied $L$ and fixed $D=16$ and $\Delta t=0.1$. (bottom) Results for varied $\Delta t$ and fixed $L=8$ and $D=16$. The different plots along each row are labeled with the time-evolution scheme used: semi-Lagrangian (SL), fourth-order Runge-Kutta (RK4), and single-site TDVP. Errors with respect to analytical results for the variance (i.e., thermal velocity squared, green), drift velocity (blue), and smoothness of the Maxwellian distribution (Eq. (17); red) are plotted in different colors. The two symbols denote calculations performed with a second-order or fourth-order centered finite-difference stencil. By construction, the stencil plays no role in SL time evolution. Missing data is a result of numerical instabilities.
@@
The results are shown in Fig. 2. Plots in the top row show that a bond dimension of $D=16$ is sufficiently large for these calculations, since the error is not reduced with increased bond dimension. In the single-site TDVP calculations, the bond dimension throughout the simulation is fixed at the bond dimension of the initial state, which is about $D=10$. Plots in the middle row show how error scales with increasing grid resolution, for fixed bond dimension $D=16$ and time step $\Delta t=0.1$. This time step is too large for RK4 with resolutions greater than $L=6$, leading to numerical instabilities. In the SL case, increasing resolution most notably improves the error in the variance (temperature) of the Maxwellian, as well of the smoothness of the resulting distribution function. For TDVP, increasing resolution also reduces errors, though at higher grid resolutions another source of error seems to appear. The bottom row compares errors for calculations with $D=16$, $L=8$ for different time steps. With SL time evolution, error in the drift velocity decreases with reduced time step, but the error in variance and smoothness increases. This suggests that the error arises from the projection of the Lagrangian dynamics onto the grid at each time step. With TDVP, for the calculation with the fourth-order finite-difference stencil, performance improves with reduced time step, as is expected since the sequential iterative updates have an associated commutation error. However, with the second-order stencil, the error appears to be dominated by lack of grid resolution since reducing time step does not reduce error.
In summary, as a local time evolution scheme, in practice, TDVP is computationally cheaper than both SL and RK4 time evolution. The above results show that both SL time evolution and TDVP allow one to take larger time steps than with conventional grid-based methods like RK4. However, note that when TDVP is performed with RK4 as the local time integrator (as is done here), there still exists a maximum time-step above which time evolution becomes unstable. In general, the error in the thermal velocity is smaller with TDVP than SL time evolution (with which the Maxwellian tends to broaden). However, the distribution function from SL tends to be smoother than the results from TDVP. This may become problematic for long-time time integration using TDVP.
@@ 6 Discussion
$T(i_{1},...i_{j},i_{j+1},...i_{L})=U(i_{1},...i_{j})\Sigma V^{\dagger}(i_{j+1},...,i_{L}),$ (46)
@@
In comparison to non-quantized low-rank methods, the proposed QTN method exhibits a much more modest reduction in computational resources for the problems discussed above. One advantage of the QTN methods is the ability to use time-steps larger than the CFL limit because of the local time evolution scheme, without having to perform semi-Lagrangian time evolution. Furthermore, QTNs are more likely to demonstrate a bigger advantage for problems requiring orders of magnitude more grid points along each dimension, as will be necessary for more realistic problems (e.g., with the proper mass ratio). A more thorough investigation on the efficiency of the QTN format for different simulation parameters needs to be performed. Additionally, testing the proposed algorithm for more complex problems, including those in different coordinate systems and especially those that are currently not tractable using conventional methods, is certainly worth pursuing. The QTN format is also not limited to a real-space, grid-based interpretation. Since the convolution operation can be performed efficiently within the QTN format (Kazeev et al., 2013b), one can also implement an efficient QTN solver with spectral bases. Spectral representations are often more compact than grid-based representations, so the degree to which QTNs can further compress these representations should be measured. Existing work also discusses finite-element solvers in the QTN format (Kazeev et al., 2017; Kazeev & Schwab, 2018; Kazeev et al., 2022; Kornev et al., 2023). Performing a careful comparison to understand the benefits and caveats of using different representations would be of practical interest and may improve our understanding of QTNs more generally.
@@ 7 Conclusion
Understanding the nonlinear behavior of collisionless plasmas requires first-principles descriptions such as provided by the coupled Vlasov-Maxwell system of equations. However, solving the Vlasov equation with traditional numerical algorithms is extremely resource-intensive, so an alternative approach that offers significantly reduced computational cost but nonetheless captures the important physical effects is very much needed. To address this problem, in this work we pioneer a grid-based Vlasov-Maxwell solver utilizing a quantum-inspired approach: the quantized tensor network (QTN) format. Moreover, we use semi-Lagrangian time integration and a QTN time evolution scheme, both of which allow for timesteps larger than the CFL limit, thus overcoming a major issue plaguing high-resolution calculations. The cost of our algorithm at worst scales roughly like $\mathcal{O}(D^{4})$, where $D$ is the rank or bond dimension of the QTN, and determines the accuracy of the calculation. In the test cases considered, we find that QTNs of at most $D=64$ are sufficient for capturing the expected dynamics with reasonable accuracy, even though the simulation is performed on grids of around $2^{36}$ grid points in total. This success suggests that the quantum-inspired representation can compress the Vlasov equation without significant loss of relevant physical effects, and thus may be a useful tool for interpreting and and exploiting the multi-scale properties of plasma dynamics. More practically, the low-rank calculation offers approximately a $10^{5}$ reduction in the number of degrees of freedom compared to the corresponding traditional finite difference representation, and a $10^{2}$ reduction with respect to a comparable finite element calculation. Due to the significant reduction in degrees of freedom and larger time-steps, these simulations require significantly fewer resources than conventional numerical methods. Critically, despite achieving much greater efficiencies, the key physical processes of the test problems investigated are retained. Therefore, we expect that QTN-based algorithms will enable Vlasov simulations that would have remained out of reach with conventional numerical methods, and thus accelerate the pace of discovery in plasma physics and associated fields.
@@ 9 Acknowledgments
EY thanks Jimmy Juno for discussions regarding the cost of the discrete-Galerkin finite element Vlasov solver in Gkeyll. The authors were supported by awards DE-SC0020264 and DE-SC0022012 from the Department of Energy.

### v4

@@ 4.3.1 Advection with time varying electric field
We perform time evolution using a second-order semi-Lagrangian (SL) scheme, fourth-order Runge-Kutta (RK4), and TDVP. The calculations use $2^{L}$ grid points along the $v_{x}$ and $v_{y}$ axes (with bounds $[-12v_{th},12v_{th})$). The test electron distribution is initialized with drift velocities $v_{x,0}=v_{y,0}=0$ and thermal velocity of 1, and is evolved until a time of 100 using time step $\Delta t$. Because split-step schemes are used, the results are shown with respect to the smallest time-step used in the time-evolution scheme ($\Delta t_{\text{step}}$) for a more fair comparison. The calculations were performed using a sequential layout with data along both axes encoded via a binary mapping. Both second-order and fourth-order centered finite difference stencils are considered.
Figure 2: Rms errors in the dynamics for advection with time-varying electric field. Test parameters are set to be $\omega=0.4567$, $E_{0}=0.9$, and $B_{0,z}=1$. Calculations are performed with bond dimension $D$, grid resolution $2^{L}$ along each dimension, and time step $\Delta t$. The finest interval of each time step is $\Delta t_{\text{step}}$. (top) Results for varied $D$ and fixed $L=8$ and $\Delta t_{\text{step}}=0.0125$. (middle) Results for varied $L$ and fixed $D=16$ and $\Delta t_{\text{step}}=0.05$. (bottom) Results for varied $\Delta t_{\text{step}}$ and fixed $L=8$ and $D=16$. The dashed black lines depict multiples of the CFL limit. The different plots along each row are labeled with the time-evolution scheme used: semi-Lagrangian (SL), fourth-order Runge-Kutta (RK4), and single-site TDVP. Errors with respect to analytical results for the variance (i.e., thermal velocity squared, green), drift velocity (blue), and smoothness of the Maxwellian distribution (Eq. (17); red) are plotted in different colors. The two symbols denote calculations performed with a second-order or fourth-order centered finite-difference stencil. By construction, the stencil plays no role in SL time evolution. Missing data is a result of numerical instabilities.
@@
The results are shown in Fig. 2. Plots in the top row show that a bond dimension of $D=16$ is sufficiently large for these calculations, since the error is not reduced with increased bond dimension. In the single-site TDVP calculations, the bond dimension throughout the simulation is fixed at the bond dimension of the initial state, which is about $D=10$. Plots in the middle row show how error scales with increasing grid resolution, for fixed bond dimension $D=16$ and time step $\Delta t_{\text{step}}=0.05$. This time step is too large for RK4 with resolutions greater than $L=7$, leading to numerical instabilities. In the SL case, increasing resolution most notably improves the error in the variance (temperature) of the Maxwellian, as well of the smoothness of the resulting distribution function. For TDVP, increasing resolution also reduces errors of the drift velocity, but the distribution function exhibits more numerical noise (larger $\epsilon_{\nabla}$) at higher grid resolutions. The bottom row compares errors for calculations with $D=16$, $L=8$ for different time steps. With SL time evolution, error in the drift velocity decreases with reduced time step, but the error in variance and smoothness increases. This suggests that the error arises from the projection of the Lagrangian dynamics onto the grid at each time step. With TDVP, for the calculation with the fourth-order finite-difference stencil, performance improves with reduced time step, as is expected since the sequential iterative updates have an associated commutation error. However, with the second-order stencil, reducing time step does not reduce error.
In summary, as a local time evolution scheme, in practice, TDVP is computationally cheaper than both SL and RK4 time evolution. The above results show that both SL time evolution and TDVP allow one to take larger time steps than with explicit grid-based methods like RK4. However, note that when TDVP is performed with RK4 as the local time integrator (as is done here), there still exists a maximum time-step above which time evolution becomes unstable. In general, the error in the thermal velocity is smaller with TDVP than SL time evolution (with which the Maxwellian tends to broaden). However, the distribution function from SL tends to be smoother than the results from TDVP. This may become problematic for long-time time integration using TDVP.
@@ 6 Discussion
$\textsf{T}(i_{1},...i_{j},i_{j+1},...i_{L})=\textsf{U}(i_{1},...i_{j})\Sigma\textsf{V}^{\dagger}(i_{j+1},...,i_{L}),$ (36)
@@
In comparison to non-quantized low-rank methods, the proposed QTN method exhibits a much more modest reduction in computational resources for the problems discussed above. One advantage of the QTN methods is the ability to use time-steps somewhat larger than the CFL limit because of the local time evolution scheme, without having to perform semi-Lagrangian time evolution. Furthermore, QTNs are more likely to demonstrate a bigger advantage for problems requiring orders of magnitude more grid points along each dimension, as will be necessary for more realistic problems (e.g., with the proper mass ratio). A more thorough investigation on the efficiency of the QTN format for different simulation parameters needs to be performed. Additionally, testing the proposed algorithm for more complex problems, including those in different coordinate systems and especially those that are currently not tractable using conventional methods, is certainly worth pursuing. The QTN format is also not limited to a real-space, grid-based interpretation. Since the convolution operation can be performed efficiently within the QTN format (Kazeev et al., 2013a), one can also implement an efficient QTN solver with spectral bases. Spectral representations are often more compact than grid-based representations, so the degree to which QTNs can further compress these representations should be measured. Existing work also discusses finite-element solvers in the QTN format (Kazeev et al., 2017; Kazeev & Schwab, 2018; Kazeev et al., 2022; Kornev et al., 2023). Performing a careful comparison to understand the benefits and caveats of using different representations would be of practical interest and may improve our understanding of QTNs more generally.
@@ 7 Conclusion
Understanding the nonlinear behavior of collisionless plasmas requires first-principles descriptions such as provided by the coupled Vlasov-Maxwell system of equations. However, solving the Vlasov equation with traditional numerical algorithms is extremely resource-intensive, so an alternative approach that offers significantly reduced computational cost but nonetheless captures the important physical effects is very much needed. To address this problem, in this work we pioneer a grid-based Vlasov-Maxwell solver utilizing a quantum-inspired approach: the quantized tensor network (QTN) format. Moreover, we use semi-Lagrangian time integration and a QTN time evolution scheme, both of which allow for timesteps somewhat larger than the CFL limit. The cost of our algorithm at worst scales roughly like $\mathcal{O}(D^{4})$, where $D$ is the rank or bond dimension of the QTN, and determines the accuracy of the calculation. In the test cases considered, we find that QTNs of at most $D=64$ are sufficient for capturing the expected dynamics with reasonable accuracy, even though the simulation is performed on grids of around $2^{36}$ grid points in total. This success suggests that the quantum-inspired representation can compress the Vlasov equation without significant loss of relevant physical effects, and thus may be a useful tool for interpreting and and exploiting the multi-scale properties of plasma dynamics. More practically, the low-rank calculation offers approximately a $10^{5}$ reduction in the number of degrees of freedom compared to the corresponding traditional finite difference representation, and a $10^{2}$ reduction with respect to a comparable finite element calculation. Due to the significant reduction in degrees of freedom and larger time-steps, these simulations require significantly fewer resources than conventional numerical methods. Critically, despite achieving much greater efficiencies, the key physical processes of the test problems investigated are retained. Therefore, we expect that QTN-based algorithms will enable Vlasov simulations that would have remained out of reach with conventional numerical methods, and thus accelerate the pace of discovery in plasma physics and associated fields.
@@ 8 Code Availability
9 Supplementary Information
Supplementary material is available at https://doi.org/10.1017/jfm.2019…
10 Declaration of Interest
The authors report no conflict of interest.
11 Funding
The authors were supported by awards DE-SC0020264 and DE-SC0022012 from the Department of Energy.
@@ 12 Acknowledgments
EY thanks Jimmy Juno for discussions regarding the cost of the discrete-Galerkin finite element Vlasov solver in Gkeyll.
13 Author ORCIDs
E. Ye, https://orcid.org/$0$-0001-9694-568X; N. F. Loureiro, https://orcid.org/$0$-0001-9755-6563.